A bus moving in a straight line at a speed of 20 m/s begins to slow at a constant rate of 3.0 m/s2. Find how far it goes before stopping.


1
Expert's answer
2021-06-21T11:39:41-0400

This distance is


"d=\frac{v^2}{2a}=\frac{20^2}{2·3}=66.7\text{ m}."